Find the tangents and normals to the curve
xy+2x-y=0
that are parallel to the line
2x+y=0
Draw the graphs showing the tangents and normal to the curve.
1
Expert's answer
2019-12-02T09:41:58-0500

Let's rewrite the equation in form

"2x=y(1-x)"

Hence

"y=\frac{2x}{1-x}"

The equation of the tangent

"y_t=y'(x_0)(x-x_0)+y(x_0)"

"y'=\frac{2(1-x)+2x}{(1-x)^2}=\frac{2}{(1-x)^2}"

Since the tangent is parallel to the line

2x+y=0 

"y'(x_0)=-2\iff \frac{2}{(1-x)^2}=-2"

The last equation has no real roots hence there are no tangents to this curve which are parallel to this line.

The equation of the normal

"y_n=-\frac{1}{y'(x_0)}(x-x_0)+y(x_0)"

Since the normal is parallel to the line

2x+y=0 

"-\frac{1}{y'(x_0)}=-2\iff -\frac{(1-x^2)}{2}=-2"

"(1-x)^2=4"

"x_1=3;\quad x_2=-1"

Two normals to the curve have the equation

"y_{n1}=-2(x-3)-3=-2x+3"


"y_{n2}=-2(x+1)-1=-2x-3"